My school wants me to simplify the expression. But I'm not sure how to I keep trying but I keep getting the wrong answer. Is there any way that you can show me in steps and explain it to so that i can learn this to complete my quiz?
-4x - 4y
________
x + y =

Answer by sudhanshu_kmr(1152)


-4x - 4y = -4(x+y)

so, -4(x+y) / (x+y) =-4
answer is: -4